# The controlled exciton transport of the Multi-chain system by cavity-dressed energy level crossings and anticrossings

### v2

@@ Abstract
The accomplished functions of a variety of quantum devices are closely associated with the controlling of exciton transport. To this end we study the exciton transport of the two-dimensional system consisting of two-level multichains with various coupling configurations in a cavity. Two types of the chains are considered, including the Tavis-Cummings and the Su-Schrieffer-Heeger chain. Two inter-chain coupling conformations are considered, including square and triangle type. The effects of the inter-chain coupling, the dimerization parameter, the cavity, the length and number of the chains on the exciton transport are in detail investigated for different coupling configurations through the spectra and steady-state dynamics of the system. The results show that in the absence of cavity the exciton transport currents and efficiency is decided by the distribution of population of exciton on the multichains. However, when the cavity is considered the exciton transport currents and efficiency is controlled by the cavity-dressed energy level crossings and anticrossings near the zero-energy modes, at which the coherent excitation and Landau-Zener transitions occur. Therefore, the exciton transport can be enhanced or suppressed at the crossings and anticrossings, in which the polariton acts as crucial role. Besides, it is discovered that the exciton transport efficiency is closely related with the parity of both the length and number of the chains. This work is important for the understanding of the exciton transport mechanism in the multichain-cavity system, and provides theoretical basis for excitonic devices with controllable and efficient exciton transport.
@@ I Introduction
In quantum device, the carriers of energy and information transport are mainly electron, photon, and exciton. Enhancing and controlling the efficiency, the speed, the distance and the current of carrier transport are the center to accomplish the functions of various quantum devices, such as organic solar cells wurfel2015impact ; lu2022simultaneously ; melianas2017charge , organic light-emitting diodes shan2011transport ; tanase2003unification ; shen2022through , quantum dot solids zhang2022ultrafast , organic semiconductor devices sneyd2021efficient , quantum simulation, and quantum computation georgescu2014quantum ; yang2019quantum etc. The exciton loss caused by different factors, including the dissipation cainelli2021exciton ; zhu2024quantum ; bose2022effect , exciton-exciton annihilation wittmann2020energy , disorder zhu2022single ; quenzel2022plasmon ; schachenmayer2015cavity and defect sites wagner2011bound ; akselrod2014visualization , is one of the influence on the efficiency in a long transport distance. Thus, a large of studies are devoted to revealing the exciton transport mechanisms and properties of various quantum system engel2007evidence ; menke2013tailored , and to achieving the requirements for the property of exciton transport through reduce and avoid detrimental factors.
@@
In experiment, the one-dimensional (1D) system of exciton or photon transport can be realized by various platforms, including organic aggregates molecule chainssu1979solitons , superconducting qubit circuits roushan2017spectroscopic ; xu2018emulating ; blais2003tunable , Rydberg-atom chains barredo2015coherent ; orioli2018relaxation , plasmons su2021perovskite ; seitz2020exciton , semiconductorssneyd2021efficient ; giannini2022exciton ; muller2023directed . At room temperature the strong exciton-photon coupling between organic aggregates and microcavities lidzey1998strong ; lidzey1999room ; kena2008strong ; tischler2005strong ; coles2014polariton or plasmons bellessa2014strong ; bellessa2004strong ; zengin2015realizing ; li2018plasmon can result in new applications like polariton lasing kena2010room , in which the molecular energy structures are modified galego2015cavity and the polaritons form between the edge state excitons and the cavity photons kavokin2003cavity ; agranovich2003cavity . The presence of polaritons can enhance the efficiency wei2019enhanced ; wei2022cavity , rate wang2021polariton and distance du2018theory ; myers2018polariton of the exciton transport because the decay of excitons on chains is suppressed.
@@
The hybrid system consisting of an 1D molecule chain and a cavity can be described by the Tavis-Cummings (TC) model tavis1968exact for uniform coupling and the Su-Schrieffer-Heeger (SSH) model su1979solitons for the staggered hopping interaction. Based on the TC model, the interaction strength of the collective $N-$atom can be predicted fink2009dressed , a new type of quantum phase transitions was discovered at quite low coupling strengthszou2013quantum . Specially, the study shows that the SSH chain can be tuned into topologically trivial and nontrivial phases for the even number of the two-level systems (TLSs) cai2019observation . For the even chain, in the topologically nontrivial regime, a small gap bewteen two localized nearly-zero-energy modes arises in the middle of the topological bandgap, which leads to the formation of the hybrid edge modes with the left and the right edge state. Therefore, the nontrivial edge states topologically protected can enhance the exciton transport for a finite-size SSH chain, which is verified by the theoretical and experimental results chang2020probing ; wei2022cavity . In fact, the nature of the gap is the energy anticrossing, at which the Landau–Zener (LZ) transitions occur. In the two-dimensional (2D) and three-dimensional tungsten disulfide system, the anticrossing dispersions were presented, at which the boosting transport capability of excitons was corroborated in theoretically and experimentally guo2022boosting ; yuan2017exciton . Thus, the properties of the more than 1D system are fascinating and valuable to study on the exciton transport.
@@
Motivated by the works on the high dimensional tungsten disulfide, we devote to the studies of the joint effect of inter-chain on exciton transport in a multi-chain TLSs with cavity in this paper. Each chain contains two alternative types, i.e., the TC and the SSH chain, and two conformations of the inter-chain coupling are set, i.e., square and triangle type. Firstly, in order to reveal the mechanism of exciton transport dynamics of the hybrid system in the condition of the strong exciton-cavity coupling, the energy spectrum with the inter-chain coupling strength and the dimerization parameter is presented through obtaining the eigenvalues of the Hamiltonian for six configurations. Meanwhile, we calculate the von Neumann entropy of exciton and Hopfield coefficients of photon to confirm the locations of the crossings and anticrossings of the near-zero energy mode dressed by the cavity. What’s more, the effects of the inter-chain coupling, the dimerization parameter, the cavity, the TLS transition frequency detuning, the number and length of the chains on the steady-state dynamics of exciton transport is investigated for different configurations by numerically solving the Lindblad quantum master equation under the TLS drive and TLS+cavity drive. In the absence of cavity, the channel of exciton transport only is the multichains, in which the efficiency depends on the distribution of exciton on the whole multichains. Importantly, for the hybrid system the exciton transport is obviously related to the crossings and anticrossings of the dressed near-zero energy modes due to the coherent excitation and LZ transition between the multichains and the cavity, at which the excitation is exchanged. As a consequence the exciton transport is controlled by the crossings and anticrossings, in which the enhancing or suppressing effects is decided by the configuration of the multichains and the polaritons act as important role. The optimal parameters regions on the space of the TLS transition frequency detuning, the inter-chain coupling strength, and the dimerization parameter are presented. Further, the results show that the exciton transport dynamics depend on the odevity of the number and length of the chains, i.e. the efficiency behaves oscillation with them.
@@
The remainder of this paper is organised as follows: In Sec. II the theoretical model is introduced including the Hamiltonian of the multichain-cavity coupling system (in Sec. II.1), the formula calculating the von Neumann entropy of excitons (in Sec. II.2), and the Lindblad quantum master equation of the system with external driving field (in Sec. II.3). In Sec. II.4 the exciton transport currents and efficiency for the steady state are derived. In Sec. III the energy spectrums are given by numerically diagonalizing the Hamiltonian. In Sec. IV we present and discuss the steady-state dynamics of excitons for various configurations of the multichains, in which the effects of the cavity, the inter-chain coupling, the dimerization parameter, the TLS transition frequency detuning (in Sec. IV.1), the number and length of the chains (in Sec. IV.2) on exciton transport are analysed. Section VI is a brief summary and future perspective.
@@
Figure 1: (color online) The schematic representation of the hybrid system that the 2D multi-chain TLSs are placed in a cavity for the TC ST (a), TC TT (b), and SSH ST (c). The red sphere represents the TLS with ground state $|g\rangle$ and excited state $|e\rangle$ between which the transition frequency is $\omega_{\rm a}$ and the decay rate is $\gamma$. The decay rate of the imperfect cavity is $\kappa$. The 2D TLSs array can be driven by an external coherent field with driving frequency $\omega$ via the first TLS of the first row with driving amplitude $\xi_{{\rm T},1}$ or/and the cavity with driving amplitude $\xi_{c}$. The excitons in the multichains finally outflow from the last column into a reservoir (drain) with rate $\gamma_{\rm R}$. For the TC chains and SSH chains, the intra-chain coupling strengths are respectively denoted by the $h_{\rm TC}$ and $h_{\rm SSH}$. $\delta$ is the dimerization strength of SSH chain. It exists the inter-chains couplings with the strength $\zeta$ between the nearest TLSs in two nearest chains.
@@ II.1 Hamiltonian
The considered system is a 2D molecular aggregate placed in a single-mode cavity, which can be modelled as multichains with $L$ ($L\geqslant 2$) chains as shown in Fig. 1. The $k$th chain has $N_{k}$ identical TLSs. For simplicity, we set that the numbers of the TLSs for different chains are the same ($N_{1}=N_{2}=...=N_{L}=N$). The Hamiltonian of the multichain-cavity coupling system can be expressed as
@@
$H_{\rm e}=\sum^{L}_{k=1}H_{{\rm e},k}+H_{\rm e,inter-chain}$ (2)
Here, $H_{{\rm e},k}$ are the Hamiltonian of the $k$th bare chain. Each chain can be set as different models.
For the uniform coupling strength between two nearest-neighbor TLS monomers in a chain, under the rotating frame of the driving frequency $\omega$ the Hamiltonian of the $k$th chain is described by the TC model as
@@
$h_{\rm TC}\sum_{i=1}^{N-1}(\sigma_{k,i}^{+}\sigma_{k,i+1}^{-}+{\rm H.c.}),$ (3)
where $h_{\rm TC}$ characterizes the coupling strength between two nearest-neighbor TLS monomers in a chain, $\sigma_{k,i}^{+}=|e_{k,i}\rangle\langle g_{k,i}|$ ($\sigma_{k,i}^{-}=|g_{k,i}\rangle\langle e_{k,i}|$) is the raising (lowering) operator of the TLS with same energy $\omega_{\rm a}$ because of the identical partical (setting $\hbar=1$) at the $i$th site of the $k$th chain with the groud state $|g_{k,i}\rangle$ and the excited state $|e_{k,i}\rangle$, $\Delta_{\rm a}=\omega-\omega_{\rm a}$ is the TLSs transition frequency detuning relative to driving field frequency $\omega$, ${\rm H.c.}$ denotes the Hermitian conjugate.
Moreover, for the staggered hopping interaction between the nearest-neighbor TLS monomers in a chain, the Hamiltonian of each chain is described by the SSH model as
$H_{{\rm e},k}=$ $-\sum_{i=1}^{N}\Delta_{\rm a}\sigma_{k,i}^{+}\sigma_{k,i}^{-}-h_{\rm SSH}\times$
$\sum_{i=1}^{N-1}\left[1+(-1)^{i}\delta\right](\sigma_{k,i}^{+}\sigma_{k,i+1}^{-}+{\rm H.c.}),$ (4)
where $h_{\rm SSH}$ is similar to the $h_{\rm TC}$, and $\delta$ represents the dimerization strength.
The inter-chain coupling term is denoted by the $H_{\rm e,inter-chain}$, whose form depends on the inter-chain coupling configuration in the 2D multichains. In this paper, we consider the two simplest types of the configurations, which are named as the square type (ST) [Fig. 1(a)] and the triangle type (TT) [Fig. 1(b,c)]. The corresponding Hamiltonian can be respectively written as
@@
where for TT the upper bounds of $L$ for odd and even term are respectively
@@
and $\zeta$ denotes the inter-chain coupling strength, namely the interaction strength between two nearest-neighbor TLSs in two nearest-neighbor chains. Therefore, the multichains have six possible configurations based on the permutation and combination of the inter-chain and the intra-chain coupling types, which are expressed as TC ST, TC TT, SSH ST, SSH TT, TC-SSH ST, and TC-SSH TT. The TC ST, TC TT, and SSH ST are schematically depicted in Fig. 1(a), Fig. 1(b), and Fig. 1(c) respectively. Correspondingly, the SSH TT is that the TC chains in the TC TT are substituted by the SSH chains. Particularly, the TC and SSH chain are staggered arrangement in TC-SSH ST and TC-SSH TT, in which we set that the odd chains are TC type and the even chains are SSH type.
@@
In order to find the energy level crossings and anticrossings of Hamiltonian (1) in the parameter space of the inter-chain coupling strength $\zeta$ and the dimerization parameter $\delta$, one can calculate the von Neumann entropy of the excitons by
@@
For the hybrid system consisting of excitons and photons, its eigenstate can be indicated as $|\psi_{\rm e-p}\rangle$. Thus the reduced density matrix $\rho_{\rm e}$ of the excitons is obtained by tracing out the degrees of freedom of the photons, i.e., $\rho_{\rm e}={\rm tr}_{\rm p}(|\psi_{\rm e-p}\rangle\langle\psi_{\rm e-p}|)$. According to the conclusion of Ref. hu2018effect , the crossings (anticrossings) correspond to the saltatorial intersections (maximum values or acclive intersections) of the von Neumann entropy with parameter. Therefore, the entanglement between upper and lower energy levels is maximal at the anticrossings, at which the LZ transitions occur. The limiting of distance between two anti-cross energy levels is crossing that means the energy degeneracy. At the crossings, the excitations are exchanged coherently between the multichains and the cavity. As a result, in energetically degenerate state the exciton transport can be observed tagarelli2023electrical and boosted lin2012efficient .

### v3

@@ Abstract
The accomplished functions of a variety of quantum devices are closely associated with the controlling of exciton transport. To this end we study the exciton transport of the two-dimensional system consisting of two-level multichains with various coupling configurations in a cavity. Two types of the chains, the homogeneous and heterogeneous coupling chain, as well as two inter-chain coupling conformations, the square and triangle type, are considered. The effects of the inter-chain coupling, the dimerization parameter, the cavity, the length and number of the chains on exciton transport are in detail investigated for different coupling configurations through the spectra, the Hopfield coefficients, and the steady-state dynamics of the system. The results show that in the absence of cavity the exciton transport currents and efficiency are decided by the distribution of exciton on the multichains. However, when the cavity is considered the exciton transport can be significantly enhanced or suppressed by the polariton at the cavity-dressed energy level crossings and anticrossings near zero-energy modes, at which the coherent excitation and Landau-Zener transitions occur. Meanwhile, we discover that the incontinuity and extreme points of the second order partial derivative of the Hopfield coefficients of the photon with the inter-chain coupling and the dimerization parameter correspond respectively to the crossings and anticrossings at the extreme points of the photon occupation number. In addition, the exciton transport efficiency is closely related to the parity of both the length and number of the chains, which behaves as oscillation. This work is important for understanding the exciton transport mechanism in the multichain-cavity system, and provides theoretical basis for excitonic devices with controllable and efficient exciton transport.
@@ I Introduction
In quantum device, the carriers of energy and information transport are mainly electron, photon, and exciton. Enhancing and controlling the efficiency, the speed, the distance, and the current of carrier transport are the center to accomplish the functions of various quantum devices, such as organic solar cells wurfel2015impact ; lu2022simultaneously ; melianas2017charge , organic light-emitting diodes shan2011transport ; tanase2003unification ; shen2022through , quantum dot solids zhang2022ultrafast , organic semiconductor devices sneyd2021efficient ; hagenmuller2017cavity , quantum simulation, and quantum computation georgescu2014quantum ; yang2019quantum etc. The factors affecting exciton transport are various, including the dissipation cainelli2021exciton ; zhu2024quantum ; bose2022effect , exciton-exciton annihilation wittmann2020energy , disorder zhu2022single ; quenzel2022plasmon ; schachenmayer2015cavity ; allard2022disorder ; balasubrahmaniyam2023enhanced ; ribeiro2022multimode , defect sites wagner2011bound ; akselrod2014visualization , and topological property buendia2024long ; allard2023multiple ; nie2021topology . Thus, a large of studies are devoted to revealing the exciton transport mechanisms and properties of various quantum system engel2007evidence ; menke2013tailored , and to achieving the requirements for the property of exciton transport through reduce and avoid detrimental factors.
@@
In experiment, the one-dimensional (1D) system carrying electron, photon, or exciton transport can be realized by various platforms, including organic aggregates molecule chainssu1979solitons , superconducting qubit circuits roushan2017spectroscopic ; xu2018emulating ; blais2003tunable , Rydberg-atom chains barredo2015coherent ; orioli2018relaxation , plasmons su2021perovskite ; seitz2020exciton , semiconductorssneyd2021efficient ; giannini2022exciton ; muller2023directed . At room temperature the strong exciton-photon coupling between organic aggregates and microcavities lidzey1998strong ; lidzey1999room ; kena2008strong ; tischler2005strong ; coles2014polariton or plasmons bellessa2014strong ; bellessa2004strong ; zengin2015realizing ; li2018plasmon can result in new applications like polariton lasing kena2010room and the polaritons forming between the edge state excitons and the cavity photons kavokin2003cavity ; agranovich2003cavity , in which the molecular energy structures are modified galego2015cavity . The presence of polaritons can adjust the efficiency wei2019enhanced ; wei2022cavity ; feist2015extraordinary , rate wang2021polariton ; aroeira2023theoretical , and distance du2018theory ; myers2018polariton ; sandik2024cavity ; takazawa2010fraction of the exciton transport because the decay of excitons on chains is suppressed.
@@
The hybrid system consisting of an 1D molecule chain and a cavity can be described by the Tavis-Cummings model tavis1968exact . The chain with the staggered hopping interaction is depicted by the Su-Schrieffer-Heeger (SSH) model su1979solitons . In the SSH chain there are topologically trivial and nontrivial phases for the even number of the two-level systems (TLSs) cai2019observation . The even chain in the topologically nontrivial regime has a small gap bewteen two localized nearly-zero-energy modes in the middle of the topological bandgap, which leads to the formation of the hybrid edge modes with the left and the right edge state. Therefore, the nontrivial edge states topologically protected can enhance the exciton transport for a finite-size SSH chain, which was verified theoretically and experimentally chang2020probing ; wei2022cavity . In fact, the nature of the gap is the energy level anticrossing, at which the Landau–Zener (LZ) transition occurs. In the two-dimensional (2D) and three-dimensional tungsten disulfide material, the anti-crossing dispersions were presented, in which the boosted transport capability of excitons was corroborated in experiment guo2022boosting ; yuan2017exciton . The properties of the multichains system coupling to a cavity were theoretically investigated kvande2023finite . Thus, the potential properties of the more than 1D system on the exciton transport are fascinating and valuable to study.
@@
Motivated by the works on the high dimensional tungsten disulfide, in this paper we devote to the studies of the inter-chain joint effect on exciton transport in 2D TLSs array with a cavity. Each chain contains two alternative coupling types, i.e., the homogeneous (HO) and the heterogeneous (HE) coupling, and two conformations of the inter-chain coupling are set, i.e., square and triangle types. Firstly, the energy spectra of the hybrid system in the condition of the strong exciton-cavity coupling with the inter-chain coupling strength and the dimerization parameter are presented by diagonalizing numerically the Hamiltonian for six configurations. Meanwhile, we calculate the von Neumann entropy of exciton to confirm the locations of the crossings and anticrossings, and the second order partial derivative of the photonic Hopfield coefficients to show the change of the occupation number of the cavity photon. What’s more, the effects of the inter-chain coupling, the dimerization parameter, the cavity, the TLS transition frequency detuning, the number and length of the chains on the steady-state dynamics of exciton transport are investigated for different configurations by numerically solving the Lindblad quantum master equation under the TLS and TLS+cavity drive. In the absence of cavity, the channel of exciton transport only is the multichains, in which its current and efficiency depend on the distribution of exciton on the whole multichains. Importantly, for the hybrid system the exciton transport is obviously related to the crossings and anticrossings of the dressed nearly-zero-energy modes due to the coherent excitation and LZ transition between the multichains and the cavity, at which the excitation is exchanged. As a consequence the exciton transport can be controlled by the crossings and anticrossings, which depend on the configurations of the multichains. Here the polaritons act as important role. In addition, the second order partial derivative of the Hopfield coefficients of the photon is a good physical quantity to measure the extreme points of the photon occupation number and the locations of the crossings and anticrossings. The optimal parameters regions on the spaces of the TLS transition frequency detuning, the inter-chain coupling strength, and the dimerization parameter are presented. Further, the results show that the exciton transport dynamics depend on the odevity of the number and length of the chains.
@@
The remainder of this paper is organised as follows: In Sec. II the theoretical model is introduced including the Hamiltonian of the multichain-cavity coupling system (in Sec. II.1), the formula calculating the von Neumann entropy of excitons (in Sec. II.2), and the Lindblad quantum master equation of the system with external driving field (in Sec. II.3). In Sec. II.4 the exciton transport currents and efficiency for the steady state are derived. In Sec. III the energy spectrums are given by numerically diagonalizing the Hamiltonian. In Sec. IV we present and discuss the steady-state dynamics of excitons for various configurations of the multichains, in which the effects of the cavity, the inter-chain coupling, the dimerization parameter, the TLS transition frequency detuning (in Sec. IV.1), the number and length of the chains (in Sec. IV.2) on exciton transport are analysed. Section V is a brief summary and future perspective.
@@
Figure 1: (color online) The schematic representation of the hybrid system that the 2D TLSs array are placed in a cavity for the three configurations HO ST (a), HO TT (b), and HE ST (c). The red sphere represents the TLS with ground state $|g\rangle$ and excited state $|e\rangle$ between which the transition frequency is $\omega_{\rm a}$ and the decay rate is $\gamma$. The decay rate of the imperfect cavity is $\kappa$. The 2D TLSs array can be driven by an external coherent field with driving frequency $\omega$ via the first TLS of the first row with driving amplitude $\xi_{{\rm T},1}$ or/and via the cavity with driving amplitude $\xi_{c}$. The excitons in the multichains finally outflow from the last column into a reservoir (drain) with rate $\gamma_{\rm R}$. The intra-chain coupling strength is $h$. $\delta$ is the dimerization strength of HE chain. It exists the inter-chains couplings with the strength $\zeta$ between the nearest TLSs in two nearest chains.
@@ II.1 Hamiltonian
The considered system is a 2D molecular aggregate placed in a single-mode cavity, which can be modelled as multichains with $L$ ($L\geqslant 2$) chains as shown in Fig. 1. The $k$th chain has $N_{k}$ identical TLSs. For simplicity, we set that the numbers of the TLSs of each chain are the same ($N_{1}=N_{2}=...=N_{L}=N$). The Hamiltonian of the multichain-cavity coupling system can be expressed as
@@
$H_{\rm e}=\sum^{L}_{k=1}H_{{\rm e},k}+H_{\rm e,inter-chain},$ (2)
where $H_{{\rm e},k}$ are the Hamiltonian of the $k$th bare chain.
Considering the staggered hopping interaction between the nearest-neighbor TLS monomers in a chain, under the rotating frame of the driving frequency $\omega$ the Hamiltonian of the $k$th chain is described by the SSH model as
@@
$\sum_{i=1}^{N-1}h\left[1+(-1)^{i}\delta\right](\sigma_{k,i}^{+}\sigma_{k,i+1}^{-}+{\rm H.c.}),$ (3)
where $h$ characterizes the hopping amplitude between two nearest-neighbor TLS monomers in a chain, $\delta$ represents the dimerization strength, $\sigma_{k,i}^{+}=|e_{k,i}\rangle\langle g_{k,i}|$ ($\sigma_{k,i}^{-}=|g_{k,i}\rangle\langle e_{k,i}|$) is the raising (lowering) operator of the TLS with same energy $\omega_{\rm a}$ (setting $\hbar=1$) at the $i$th site of the $k$th chain with the groud state $|g_{k,i}\rangle$ and the excited state $|e_{k,i}\rangle$ because of the identical partical, $\Delta_{\rm a}=\omega-\omega_{\rm a}$ is the TLSs transition frequency detuning relative to driving field frequency $\omega$, ${\rm H.c.}$ denotes the Hermitian conjugate. For $\delta=0$ and $\delta\neq 0$, the SSH chain behaves entirely different property, which correspond respectively homogeneous and heterogeneous coupling in a molecular chain. In addition, it exist the topologically nontrivial regime for $\delta\neq 0$. In order to emphasize the difference of the exciton transport of the two coupling cases, we designated separately them as the homogeneous (HO) and the heterogeneous (HE) coupling chain in the next content.
The inter-chain coupling term is denoted by the $H_{\rm e,inter-chain}$, whose form depends on the inter-chain coupling configuration in the 2D multichains. In this paper, we consider the two simplest types of the configuration, which are named as the square type (ST) [Fig. 1(a)] and the triangle type (TT) [Fig. 1(b,c)]. The corresponding Hamiltonian can be respectively written as
@@
where for TT the upper bounds of $L$ in odd and even term are respectively
@@
and $\zeta$ denotes the inter-chain coupling strength, namely the interaction strength between two nearest-neighbor TLSs in two nearest-neighbor chains. Therefore, the multichains have six possible configurations based on the permutation and combination of the inter-chain and the intra-chain coupling types, which are expressed as HO ST, HO TT, HE ST, HE TT, HOCC-HE ST, and HO-HE TT. The HO ST, HO TT, and HE ST are schematically depicted in Fig. 1(a), Fig. 1(b), and Fig. 1(c) respectively. Correspondingly, the HE TT is that the HO chains in the HO TT are substituted by the HE chains. Particularly, the HO and HE chain are arranged with staggered form in HO-HE ST and HO-HE TT, in which we set that the odd chains are HO and the even chains are HE.
@@
In order to find and distinguish accurately the locations of the energy level crossings and anticrossings of Hamiltonian (1) in the parameter space of the inter-chain coupling strength $\zeta$ and the dimerization parameter $\delta$, one can calculate the von Neumann entropy of the excitons by
@@
For the hybrid system consisting of excitons and photons, its eigenstates can be indicated as $|\psi_{\rm e-p}\rangle$. Thus the reduced density matrix $\rho_{\rm e}$ of the excitons is obtained by tracing out the degrees of freedom of the photons, i.e., $\rho_{\rm e}={\rm tr}_{\rm p}(|\psi_{\rm e-p}\rangle\langle\psi_{\rm e-p}|)$. According to the conclusion of Ref. hu2018effect , the crossings (anticrossings) correspond to the saltatorial intersections (maximum values or acclive intersections) of the von Neumann entropy with parameter. Therefore, the entanglement between upper and lower energy levels is maximal at the anticrossings, at which the LZ transitions occur. The limiting of distance between two anti-cross energy levels is crossing that means the energy degeneracy. At the crossings, the excitations are exchanged coherently between the multichains and the cavity. As a result, in energetically degenerate state the exciton transport can be observed tagarelli2023electrical and boosted lin2012efficient .